In a transistor, beta=45, the voltage across 5k? resistance which is connected in the collector circuit is 5volts. Find the base current.


1
Expert's answer
2022-02-23T12:11:45-0500

Solution

The collector current is

"I=\frac{V}{R}\\=\frac{5}{5000}\\=0.001A"

The current gain for transistor is

"I_c=\frac{I}{\beta}\\=\frac{0.001}{45}\\=2.22*10^{-5}A"